# Automated Black-box Prompt Engineering for Personalized Text-to-Image Generation

### v1

@@ 4.4 Ablation Study
4.4.1 Comparison with GPT-4V.
PRISM  relies on having a strong multimodal LLM. In our case, we chose GPT-4V as the multimodal foundation model. While in principle we may use any multimodal LLM, it is nonetheless useful to understand what benefits PRISM  adds to an already capable foundation model. To show the effectiveness of iterative prompt refinement and parallel search, we compare our method with GPT-4V’s zero-shot performance with the same system prompts for object and image inversion tasks on SDXL-Turbo. We see in Table 7 that PRISM  consistently outperforms GPT-4V’s zero-shot performance, although the latter is already compelling. In Figure 7, we show some examples of the generated results. We see that qualitatively GPT-4V can capture the high-level semantics of the reference images but still misses more fine-grained details.
4.4.2 Trade-off between N and K
@@
PRISM has two hyperparameters $N$ and $K$ which control the amount of parallel search and the depth of iterative refinement. Figure 9 shows a trade-off between N and K with the same budget $N\times K=30$. Similar to the findings of [3], we find that performance can degrade if the refinement is repeated too many times (i.e., $K$ is too large), and in general, we do not recommend practitioners with small budgets to go beyond $K=5$. Unlike jailbreaking [3], we observe that the optimal $N$ and $K$ can vary depending on the task: if the target concept is simple (e.g. a commonly seen dog), then small $N$ and $K$ are generally sufficient, and prioritizing $N$ tends to be more helpful. However, if the target concept is rarer and more complicated (e.g. a very specific toy), a larger reasoning depth (i.e., larger $K$) would be more helpful. In Figure 9, we show the distribution of iteration numbers at which the best prompt is found in the image inversion experiment. In practice, one may tune these hyperparameters further for specific use cases.
4.5 Prompt Editing
Figure 10: Prompt editing demonstration with Midjourney.
Because the prompts produced by PRISM is very human-interpretable, after obtaining a prompt from the reference images, one can easily modify the output prompts to change attributes in their desired generated images. Figure 10 demonstrates an examples of prompt editing with PRISM in Midjourney. With simple and intuitive prompt edit, we are able to change specific attributes of the images while keeping the other components in the scene unchanged.
5 Conclusion and Broader Impact Statement
In this paper, we propose PRISM, an algorithm that automatically creates human-interpretable and accurate text prompts for text-to-image generative models, based on visual concepts provided by reference images. Our method iteratively refines the sampling distribution of the text prompt via LLM in-context learning and is capable of creating prompts that are transferable to any T2I model, including black-box platforms like Dall-E and Midjourney. However, just as LLMs are suceptible to being jailbroken or adversarially manipulated by malicious actors [48], our method may also be vulnerable to malicious intent, potential bias, or limitations in the base models. Therefore, we intent to implement necessary safeguards upon the public release of our code and are committed to keep up with future advancements in improving the safety of our method.
Acknowledgements
This work was supported in part of the ONR grant N000142312368. We also thank Sony AI for sponsoring part of this work. A.R., H.H., and G.P. are supported by the NSF Institute for CORE Emerging Methods in Data Science (EnCORE). A.R. is also supported by an ASSET Amazon AWS Trustworthy AI Fellowship. Y.J. is supported by Google PhD Fellowship.
Appendix 0.A Additional Experiment Details
In this section, we provide further details about the implementation of our experiments. For all quantitative analysis that uses Stable Diffusion based model, we generate four images for each combination of prefixes and prompts. For all experiments with Dall-E based model, we generate one image per combination. In the DreamBooth dataset experiment, we also replace the class noun for “stuffed animal” with “toy” to obtain fair comparisons with textual inversion, which can only take a single token as the initialization token.
When evaluating the judge scores $\mathbf{D}(x,\hat{x})$ in PRISM iterations, we shuffle the reference images when $M>1$. The judge score is rescaled into a range from 0 to 10. During re-evaluation, we re-evaluate each prompt once with each reference image when $M>1$ and re-evaluate each prompt twice and tally these scores with the in-iteration scores when $M=1$.
Appendix 0.B Designing System Prompts
System prompting is the standard way to condition a general purpose LLM for specific tasks of request. The key idea is that, before the conversation starts, the LLM receives a tailored message, the system prompt, that provides the contexts, conversation scenario settings, formats and other guidelines as the prefix of the entire conversation ahead. In this section, we elaborate on the design of the system prompts for the prompt engineer assistant ${\mathbf{F}}$ and the judge ${\mathbf{D}}$. We also provide the full system prompts used in all of our experiments at the end of this paper in Section 0.F and in our demo code.
0.B.1 Prompt Engineer Assistant ${\mathbf{F}}$
To design the system prompts for the prompt engineer assistant ${\mathbf{F}}$, we follow [3] and include the following components in the system prompt of ${\mathbf{F}}$.
0.B.1.1 Setting
We first set up the scenarios and assign a role for the LLM to better perform on the specific task of choice. The setting paragraphs start with “You are a helpful prompt engineer assistant. You are free to generate sentences that do NOT follow English grammar. You must obey all of the following instructions.” and continue with the specific description of the task and the objective. We also inform the assistant that it is expected to iterate and refine the prompts it generates throughout the conversation.
0.B.1.2 Format
We then provide the guidelines for formatting the inputs and the outputs of the assistant. We describe what are expected in the inputs at each iteration and the content required in the outputs. We also provide descriptions of the meanings of each input and output components. More specifically, we inform the assistant that the inputs consist of three parts: a generated image, a reference images and a visual similarity score, and that the assistant is expected to generate both the improvement to refine the previous prompt and the next new prompt. All generated text are formatted in JSON.
0.B.1.3 Examples
0.B.2 Judge ${\mathbf{D}}$
Appendix 0.C Additional Results
0.C.1 Additional Qualitative Results
In Figure 13, 14, 15, 16 and 17, we provide additional qualitative showcases for subject-driven personalized T2I generation, style-driven personalized T2I generation, direct image inversion and prompt editing. We also provide an example of the iteration and refinement process as a conversation between all three components in PRISM in Figure 18.
0.C.2 Flexible Model Choices
To further demonstrate the effectiveness and flexibility of PRISM, we also experiment a different T2I Generator ${\mathbf{G}}$ and showcase the transferability of the prompts generated by PRISM. Figure 19 shows qualitative examples of PRISM prompts with Dall-E 2 as the Generator ${\mathbf{G}}$ for personalized T2I generation and the images generated from those prompts using SDXL-Turbo, Dall-E 3 and Midjourney. Our method is capable of producing human-interpretable and accurate prompts for both subject-driven T2I personalization and style-driven T2I personalization with this new Generator ${\mathbf{G}}$.
Appendix 0.D Additional Ablation Study
In this section, we provide a more detailed ablation study on each component of the PRISM framework. In particular, we demonstrate the effect of the existence of the Judge ${\mathbf{D}}$ and re-evaluation, different choices of the total budget, number of streams $N$ and number of iterations $K$, and also compare a non-LLM judge (a CLIP judge) against our choice of a LLM judge (GPT-4V Judge).
Next we take a closer look at the effect of increasing the total budget in PRISM in small budget settings. Figure 12 and 12 show the effect of increasing the number of streams $N$ and the number of iterations respectively. We observe that when increasing the number of streams $N$ while keeping the number of iterations $K$ fixed, we can obtain steady performance improvements in both human readability and prompt accuracy. When increasing $K$ while keeping $N$ fixed, although we do not observe a monotonic relationship between the performance and $K$, we can still notice a general upward trend in prompt accuracy. Generally speaking, the optimal number of iterations vary case to case, and we encourage practitioners to experiment with different choices of $K$ to obtain the best resutls.
@@ Appendix 0.D Additional Ablation Study
Finally, we demonstrate the importance of using a multimodal LLM as the Judge. When assessing image similarity, it is natural to default to existing metrics that do not involve LLM’s such as CLIP similarity. However, as we have mentioned in the main text, these metrics do not perform well outside of their trained notion of similarities and therefore is not very generalizable to custom tasks from users. Figure 21 demonstrates the qualitative difference between PRISM with a CLIP judge versus PRISM with a GPT-4V judge. We can observe that in subject-driven T2I personalization, CLIP judged PRISM often include irrelevant elements such as the environment (e.g. “on green grass”) and omits important details such as the color and the other distinctive features whereas GPT-4V judged PRISM can adhere better to object oriented details and ignores other unrelated factors. In style-driven T2I personalization, CLIP judged PRISM fails to capture the artistic styles and mainly focus on the general contents of the reference image. On the contrary, GPT-4V judged PRISM produces much more precise and focused prompts for the reference styles.
@@
Figure 11: Ablation study on different numbers of streams $N$ with the same number of iterations $K=5$.
Figure 12: Ablation study on different numbers of streams $K$ with the same number of iterations $N=3$.
Appendix 0.E Limitations and Future Works
Firstly, as we can observe in almost all of the qualitative examples, when the targeting concept is more challenging (e.g. a very particular toy), our method still fail to capture all the fine grained details in the image generation. Although this phenomenon is to some extent expected due to the fact that text-to-image generation is not a one-to-one function, there is still a long way to go in order to achieve the same performance as methods like DreamBooth [26] that involve finetuning. Moreover, even with very accurate prompts, because of the limitation of the downstream testing T2I models, sometimes it still fail to generate the correct concepts. One potential direction is to combine gradient-based search methods like PEZ [39] with PRISM to create model-specific prompts similar to CLIP-Interrogator.
@@ Appendix 0.E Limitations and Future Works
Appendix 0.F Full System Prompts
0.F.1 Subject-Driven Personalized T2I Generation
0.F.1.1 Prompt Engineer Assistant ${\mathbf{F}}$
@@ 0.F.1.2 Judge ${\mathbf{D}}$
0.F.2 Style-Driven Personalized T2I Generation
0.F.2.1 Prompt Engineer Assistant ${\mathbf{F}}$
@@ 0.F.2.2 Judge ${\mathbf{D}}$
0.F.3 Direct Image Inversion
0.F.3.1 Prompt Engineer Assistant ${\mathbf{F}}$
@@ 0.F.3.2 Judge ${\mathbf{D}}$
Figure 16: Qualitative examples of the direct image inversion task.

### v2

@@ Appendix D Additional Ablation Study
Trade-off between N and K
@@
PRISM has two hyperparameters $N$ and $K$ which control the amount of parallel search and the depth of iterative refinement. Figure 20 shows a trade-off between N and K with the same budget $N\times K=30$. Similar to the findings of Chao et al. (2023), we find that performance can degrade if the refinement is repeated too many times (i.e., $K$ is too large), and in general, we do not recommend practitioners with small budgets to go beyond $K=5$. Unlike jailbreaking (Chao et al., 2023), we observe that the optimal $N$ and $K$ can vary depending on the task: if the target concept is simple (e.g. a commonly seen dog), then small $N$ and $K$ are generally sufficient, and prioritizing $N$ tends to be more helpful. However, if the target concept is rarer and more complicated (e.g. a very specific toy), a larger reasoning depth (i.e., larger $K$) would be more helpful. In Figure 20, we show the distribution of iteration numbers at which the best prompt is found in the image inversion experiment. In practice, one may tune these hyperparameters further for specific use cases.
Comparison between a CLIP Judge and a VLM Judge
@@
Finally, we demonstrate the importance of using a VLM as the Judge. When assessing image similarity, it is natural to default to existing metrics that do not involve LLM’s such as CLIP similarity. However, as we have mentioned in the main text, these metrics do not perform well outside of their trained notion of similarities and therefore is not very generalizable to custom tasks from users. Figure 21 demonstrates the qualitative difference between PRISM with a CLIP judge versus PRISM with a GPT-4V judge. We can observe that in subject-driven T2I personalization, CLIP judged PRISM often include irrelevant elements such as the environment (e.g. “on green grass”) and omits important details such as the color and the other distinctive features whereas GPT-4V judged PRISM can adhere better to object oriented details and ignores other unrelated factors. In style-driven T2I personalization, CLIP judged PRISM fails to capture the artistic styles and mainly focus on the general contents of the reference image. On the contrary, GPT-4V judged PRISM produces much more precise and focused prompts for the reference styles. The drawbacks of using CLIP-based Judge can potentially attribute to its incapability of identifying fine grained details and distinctions in different contexts, as studied in (Thrush et al., 2022; Yuksekgonul et al., 2022). Using an autoregressively trained VLM such as GPT-4V can mitigate this issue. However, these models are not perfect either. As future works, we can potentially introduce more rule-based reasoning in the iterative process similar to (Mañas et al., 2024).
@@
Effect of the Judge ${\mathbf{D}}$ and Re-Evaluation
Prompt Length
We use the default prompt length for all baselines, which in most cases corresponds to their optimal length. For many of these methods, increasing the prompt length does not necessarily lead to better performance. For instance, in PEZ’s Section 4.2 “Prompt Length” paragraph, they explicitly note that a length of 16—rather than the longest tested length—yields the best generalizability. To further eliminate the possibility of unfair comparisons, we have also conducted an additional ablation study with GPT-4o-mini on the effect of prompt length for our model.
Table 6 we demonstrate the quantitative comparison between our method with various prompt length and PEZ, the baseline method constrained by prompt length, with its optimal length. Our results show that while PRISM benefits from longer prompt lengths, it consistently maintains high performance even with shorter prompts and significantly outperforms PEZ. Notably, we observed that as constraints on prompt length increase, PRISM tends to deviate from conventional coherent English sentences, similar to strategies employed by human prompt engineers. Additionally, unlike discrete optimization methods, longer prompt lengths do not pose significant challenges to the optimization problem inherent to our approach. It is also important to emphasize that all our experiments were conducted under the constraint that no prompt exceeds the maximum length accepted by the target T2I model, and any prompts exceeding this limit were appropriately chunked.
Table 6: Ablation study on prompt length in comparison to baseline PEZ.
Method NLL $\downarrow$ CLIP-I $\uparrow$ DINO $\uparrow$
PEZ Token Length 16 6.188 0.722 0.418
PRISM Token Length 16 4.593 0.745 0.462
PRISM Token Length 32 4.043 0.744 0.482
PRISM 3.498 0.768 0.493
Judge score distribution Comparison between the First Iteration and the Final Prompts
We also include a judge score distribution comparison between the first iteration and the final prompts. As we can observe from Figure 24 and Figure 24, in the first iteration, the most common scores obtained are 0 and 1, whereas the final prompts obtain score 7 and 8 the most. This suggests a significant improvement of the prompt quality and effectiveness throughout the iterative process. Additionally, we would like to note that as we have mentioned above and in Section 3.3, grammatical correctness is not always indicative of effective prompts and as we can observe from the qualitative examples (Figure 22), as long as NLL reaches below 3.5, further lower NLL does not result in a qualitatively noticeable difference in terms of prompt quality.
Figure 23: The Judge score distribution in the first iteration for the image inversion experiment.
Figure 24: The Judge score distribution of the final prompts for the image inversion experiment.
Appendix E Limitations and Future Works
@@
Firstly, as we can observe in almost all of the qualitative examples, when the targeting concept is more challenging (e.g. a very particular toy), our method still fail to capture all the fine grained details in the image generation. Although this phenomenon is to some extent expected due to the fact that text-to-image generation is not a one-to-one function, there is still a long way to go in order to achieve the same performance as methods like DreamBooth (Ruiz et al., 2023) that involve finetuning. In addition, one potential root of this issue can be related to VLM’s incapability to properly identify compositionality, similar to some challenges pointed out by Thrush et al. (2022); Yuksekgonul et al. (2022). Moreover, even with very accurate prompts, because of the limitation of the downstream testing T2I models, sometimes it still fail to generate the correct concepts. One potential direction is to combine gradient-based search methods like PEZ (Wen et al., 2023) with PRISM to create model-specific prompts similar to CLIP-Interrogator.
A potential concern with our method is the financial cost, as our best performing results use paid models like GPT-4. While this is valid, it’s important to note that the cost of closed sourced high-performance models has already significantly decreased. As demonstrated in Section C.2, PRISM’s performance only has marginal difference when switching from GPT-4V to GPT-4o-mini, yet GPT-4V costs \$10 per 1M input tokens and \$30 per 1M output tokens, whereas GPT-4o-mini costs only \$0.15 and \$0.6 respectively (as of October 1st, 2024). We expect the costs of these advanced models to further decrease in the future. In addition, given the rapid improvements in open-source models, we are optimistic that models like IDEFICS2 can eventually rival GPT-4V. Furthermore, generated prompts for specific objects, styles, or other visual concepts can be saved and reused for future tasks and multiple T2I platforms, not just a single generation. PRISM’s flexible cost management allows for tailored computational budgets as the choice of $N$ and $K$ can be adjusted based on specific financial and computational needs.
In addition to financial cost, our method also requires longer inference time for its best performance. Table 8, we report the latency comparison along side with other qualitative metrics on a single NVIDIA A6000 GPU. While it is true that our method may run slower when using a less efficient VLM, it also allows for flexibility in budget (the choices of $N$ and $K$ as demonstrated in Section 4.4) and model selection, in order to reduce latency while still achieving competitive performance.
Table 8: Latency comparison between our method and the baselines on the task of Dreambooth personalization on SDXL-Turbo. All PRISM variations have budget $N\times K=40$.
Method NLL $\downarrow$ CLIP-I $\uparrow$ DINO $\uparrow$ Time (s) $\downarrow$
Textual Inversion (SDXL) - 0.771 0.504 1979.303
CLIP-Interrogator 4.361 0.756 0.490 41.106
BLIP2 4.378 0.729 0.456 1.650
PEZ 6.188 0.722 0.418 179.576
PRISM (IDEFICS2) 3.047 0.739 0.468 224.451
PRISM (GPT-4o-mini) 3.498 0.768 0.493 677.076
PRISM (GPT-4V) 3.466 0.770 0.499 914.479
Appendix F Additional Related Works on LLMs as Optimizers
In this section, we would like to extend the discussion on related works on LLMs as optimizers in the current literature.
Several methods have applied the techniques of LLMs as optimizers to various vision-language downstream tasks. In particular, Mañas et al. (2024) leverage a rule-based algorithm to improve prompt-image alignment using LLM refinements, without leveraging any reference images. Hao et al. (2024) performs the same task but with a fine-tuned LLM. Liu et al. (2024) addresses traditional distriminative tasks such as image classification using a similar approach.
The most related work to our method is Idea2Img (Yang et al., 2023), but it focuses on generating a single best image rather than a generalizable prompt. In other words, Idea2Img only outputs a single best image tailored to a specific T2I model, prioritizing image quality for that one specific image without concern for the generalizability of the resulting prompts. In contrast, our method targets the generation of a generalizable prompt that works across different random seeds, contexts, and T2I platforms. This distinction accounts for the stochasticity in T2I models, where prompts must consistently produce high-quality outputs rather than relying on one best-case scenario. Unlike Idea2Img, which narrows its focus by selecting the best image at each iteration and outputs only a final image, we maintain independent streams throughout the process and use re-evaluation to identify the most effective prompt. Our approach enables broader applicability and ensures the prompts are robust and versatile across diverse scenarios.
To highlight the differences between PRISM and Idea2Img, we modify Idea2Img to output prompts and tested both methods on the DreamBooth task using SDXL-Turbo as the target and testing T2I model. Since Idea2Img only outputs an image, it is not naturally applicable to our tasks. As a result, to ensure its applicability, we modify Idea2Img to output the prompt that produces Idea2Img’s output image.
Table 9 demonstrates the quantitative comparison between PRISM and Idea2Img. PRISM outperforms Idea2Img in most metrics, particularly in generalizability (CLIP-T), which measures contextual flexibility. Qualitative comparisons in Figure 25 further demonstrate that PRISM generates prompts with greater detail and contextual relevance, avoiding irrelevant or omitted information often seen in Idea2Img’s outputs. While Idea2Img achieves lower NLL, we note (as discussed in Section 3.3) that grammatical correctness is not always indicative of effective prompts and therefore fully coherent English sentences are not always the most effective prompts. Overall, both qualitative and quantitative comparisons show that PRISM strikes a better balance between human interpretability and prompt accuracy.
Table 9: Quantitative comparison between PRISM and Idea2Img on the DreamBooth personalization task with GPT-4o-mini as the VLM backbones and SDXL-Turbo as both the target T2I model and the testing T2I model.
Method NLL $\downarrow$ CLIP-I $\uparrow$ DINO $\uparrow$ CLIP-T $\uparrow$
Idea2Img 2.657 0.759 0.485 0.219
PRISM 3.498 0.768 0.493 0.233
Figure 25: Qualitatibe comparison between PRISM and Idea2Img.
In Liu et al. (2024), they have also applied their algorithm, which is designed for discriminative tasks such as image classification, to the image inversion task. However, there are significant differences between Liu et al. (2024) and our paper in terms of algorithmic design. In particular, Liu et al. (2024) did not include the following components in their algorithm: (1) In Liu et al. (2024), they do not instruct the VLM to produce chain-of-thought improvements. In fact, in their official implementation, they specifically prompt the VLM to “Respond only with the revised text prompt and exclude any additional commentary”. (2) Liu et al. (2024) does not incorporate an external judge model to provide signals for the iterative improvements. (3) Because of the lack of a judge model, Liu et al. (2024) is unable to perform re-evaluation. Both the judge model and re-evaluation have been proven crucial for our algorithm in the ablation study we conduct in Section D in the appendix. (4) Because of the lack of re-evaluation, Liu et al. (2024) is unable to perform parallel search since there is no way for them to identify the best prompts from the search. In fact, they can only assume that VLM can monotonically improve the prompt throughout the iterations, which we have proven to be not true in our ablation study. In the case of image classification, which is the main focus of their paper, they have described an alternative way to perform this search by leveraging the validation set and the classification error. However, with the image generation task, they were not able to find a straightforward way to incorporate these designs. As a matter of fact, they use $n_{restart}=1,n_{reset}=1,m=1$ in their official implementation, which effectively makes this algorithm into re-prompting the VLM for several iterations in a single stream, without parallel search or beam search.
To demonstrate the importance of these algorithmic differences, we have tested Liu et al. (2024) in our image inversion task with GPT-4V and SDXL-Turbo. To ensure fair comparison, we have also included PRISM with the same iteration budget without parallel search (N = 1, K = 5) and the GPT-4V zero-shot results. Table 10 and Figure 26 are the quantitative and qualitative comparison between our method and Liu et al. (2024).
As we can observe, not only does Liu et al. (2024) underperform both PRISM versions, it even underperforms GPT-4V zero-shot with the system prompts we design for PRISM. This experiment shows the effectiveness of all components in our algorithm that Liu et al. (2024) misses.
Table 10: Quantitative comparison between PRISM, GPT-4V Zero Shot, and Liu et al. (2024) in image inversion task tested on SDXL-Turbo.
Method NLL CLIP-I
Liu et al. (2024) 2.520 0.720
GPT-4V Zero Shot 2.356 0.756
PRISM (N=1, K=5) 2.809 0.770
PRISM 2.762 0.776
Figure 26: Quantitative comparison between PRISM and Liu et al. (2024) in image inversion task tested on SDXL-Turbo.
Appendix G Full System Prompts
G.1 Subject-Driven Personalized T2I Generation
G.1.1 Prompt Engineer Assistant ${\mathbf{F}}$
@@ G.1.2 Judge ${\mathbf{D}}$
G.2 Style-Driven Personalized T2I Generation
G.2.1 Prompt Engineer Assistant ${\mathbf{F}}$
@@ G.2.2 Judge ${\mathbf{D}}$
G.3 Direct Image Inversion
G.3.1 Prompt Engineer Assistant ${\mathbf{F}}$